Ecker Company reports $1,800,000 of net income and declares $252,000 of cash dividends on its preferred stock for the year. At year-end, the company had 260,000 weighted-average shares of common stock.

1. What amount of net income is available to common stockholders?

What is the company’s basic EPS?
Answer #1

net income available to common stockholders=Net income-Cash dividends to preferred stock

=(1,800,000-252,000)=$1548000

Basic EPS=(Net income-Cash dividends to preferred stock)/weighted-average shares of common stock

=(1,800,000-252,000)/260,000

=$5.95(Approx).